How many molecules are in 2.4x10-26 moles of iron (II) oxide?


1
Expert's answer
2021-01-22T06:24:54-0500

1 mole = 6.023×1023 molecules

2.4×10-26 mole= 2.4×10-26 (6.023×1023)

= 14.4552 × 10-3